OpenStudy (curry):

solve for x in the interval [o,2pi) cos^2+sin +1

OpenStudy (curry):

thiss one is harder and cos4x-cos2x=0

OpenStudy (bahrom7893):

When in doubt, turn to wolf lol: http://www.wolframalpha.com/input/?i=solve%28cos4x-cos2x%3D0%29

OpenStudy (curry):

im not allowed to use a calculator

OpenStudy (bahrom7893):

i honestly don't know how to solve this without one, but likeL Cos4x-Cos2x=0 is easy to see: Cos4x=Cos2x so x=0, cuz Cos0 is 1

OpenStudy (mertsj):

Are we talking about cos^2xsinx+1?

OpenStudy (curry):

well that one but the co2x-cos2x i much harder so that too

OpenStudy (mertsj):

cos(2x)-cos(2x)=0

OpenStudy (curry):

wait not 2x it cos(4x)-cos(2x)=0

OpenStudy (mertsj):

Let me work on it for a minute.

OpenStudy (mertsj):

Let 2x=y \[\cos 2y-\cos y=0\]

OpenStudy (mertsj):

\[\cos 2y=2\cos ^2y-1\]

OpenStudy (curry):

that what u poted is another problem

OpenStudy (mertsj):

\[2\cos ^2y-1-\cos y=0\]

OpenStudy (mertsj):

\[2\cos ^2y-\cos y-1=0\]

OpenStudy (mertsj):

\[(2\cos y+1)(\cos y-1)=0\]

OpenStudy (mertsj):

\[\cos y=-\frac{1}{2}, \cos y=1\]

OpenStudy (mertsj):

\[y=\frac{2\pi}{3}, \frac{4\pi}{3},0\]

OpenStudy (mertsj):

But 2x=y so \[2x=\frac{2\pi}{3}, x=\frac{\pi}{3}\]

OpenStudy (curry):

wait i dont get hgow u went from cos2y-cosy t cos2y=2cos^2y-1 i mean u get that it is one of the indetities but sshould-1 becausse 2x= y

OpenStudy (mertsj):

\[2x=\frac{4\pi}{3}, x=\frac{2\pi}{3}\]

OpenStudy (curry):

but houldn't it be cos2y=2coss^2x-1

OpenStudy (mertsj):